OpenStudy (anonymous):

Graph the line that passes through the given point and has the given slope. (Objective 3) (3, 1) m = 1/3

OpenStudy (anonymous):

ok, do you know how to do this a little bit.

OpenStudy (anonymous):

not at all please help quickly, i don't have any time

OpenStudy (anonymous):

les than an hour with over 10 hard questions left

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

this is the formula y=mx+b

OpenStudy (anonymous):

then y=mx+b (3, 1) m = 1/3 if m=1/3 that means you plug it in and b=the y intercept or point or nuber witch is 1 and x is 3. y=1/3(3)+1

OpenStudy (anonymous):

\[y=1/3(3)+1\]

OpenStudy (anonymous):

so y=2 ?

OpenStudy (anonymous):

y=2

OpenStudy (anonymous):

CORRECT!!!!!!!!!!!!!!!!!!!!!!!!!! :), nice job

OpenStudy (anonymous):

you should create your own problem like this and do it over and over again so you may understand it even more better.

OpenStudy (anonymous):

can you do the other 10 too, if you need help tap me

OpenStudy (anonymous):

ok... so i'm still confuse as to how i graph the line?

OpenStudy (anonymous):

what do I graph?

OpenStudy (anonymous):

and thank you

OpenStudy (anonymous):

ao, i am sorry ok how to graph a line lets see

OpenStudy (anonymous):

let say you have the same problem like you had before, the question tell you. Graph the line Graph the line that passes through the given point and has the given slope. (Objective 3) (3, 1) m = 1/3

OpenStudy (anonymous):

We know that y=2 and you will do y=mx+b and

OpenStudy (anonymous):

and what?

OpenStudy (anonymous):

2 = 1/3x + b ?

OpenStudy (anonymous):

please help i have 10 minutes left!!!!!!!

OpenStudy (anonymous):

ao ok

OpenStudy (anonymous):

y=mx+b y=m3+1/1

OpenStudy (anonymous):

m3 or 3x?

OpenStudy (anonymous):

3x

OpenStudy (anonymous):

and 1/1 = 1 right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

i will be right there with you just give me a minute

OpenStudy (anonymous):

that is the graph for your equation.

OpenStudy (anonymous):

time passed already.

OpenStudy (anonymous):

ao did you see what i posted on

OpenStudy (anonymous):

let me give you an example to on how to do a graph

OpenStudy (anonymous):

we have an equation right eg. y=2x+3 then you put it like this y=2x/1+3

OpenStudy (anonymous):